# EC-DiT: Scaling Diffusion Transformers with Adaptive Expert-Choice Routing

### v2

@@ 3.5 Comparison with Token-choice Routing
Figure 5 compares EC-DiT with models scaled using token-choice MoE, as proposed in Lepikhin et al. (2020a). This routing strategy has been widely adopted in several sparse diffusion models (Yatharth Gupta, 2024; Xue et al., 2023; Fei et al., 2024). The token-choice baseline uses top-2 routing, which matches the activation size of EC-DiT with a capacity factor of $f_{c}=2.0$. Our method consistently demonstrates superior training convergence and performance throughout the entire training period. Notably, EC-DiT with 8 experts rivals the token-choice baseline with 16 experts in both generation quality and text-image alignment, while EC-DiT with more experts significantly outperforms this token-choice baseline. One reason for the slower training convergence in token-choice routing is the equal compute allocation to different image areas, which makes the compute less adaptive. Detailed image areas may not receive sufficient computation to render complex patterns, while simpler areas waste the computation they receive. Moreover, token-choice routing processes each token independently without considering sequence information, leading to a lack of perception of the overall relationships between local tokens. Additionally, the inherent load imbalance in token-choice routing results in some experts being overloaded with an excessive number of tokens and creating a training bottleneck (Zhou et al., 2022). Appendix B provides a comprehensive comparison of EC-DiT and the token-choice baseline regarding GenEval performance. As shown in Table 4, EC-DiT outperforms the token-choice model across all tested configurations, which further validates the advantages of EC-DiT in text-image alignment.
@@ 4 Related Work
Many MoE models utilize a token-choice routing strategy, where each token independently selects the top-$k$ experts for processing. Sparsely Gated MoE (Shazeer et al., 2017) first introduced top-$k$ gating in LSTM models. Such routing strategy was later extended to transformers for language modeling in GShard (Lepikhin et al., 2020a), Switch Transformer (Fedus et al., 2021), GLaM (Du et al., 2022), DeepSpeed-MoE (deepspeed), and ST-MoE (Zoph et al., 2022). To address potential load imbalances among experts, these token-choice approaches typically incorporate an auxiliary load-balancing loss (Shazeer et al., 2017). Alternatively, Zhou et al. (2022) proposed an expert-choice routing in text domain that selects the most suitable text tokens for each expert. Inspired by this, we adapt and scale up the DiT model using expert-choice routing tailored specifically for diffusion-based generative tasks. With this approach, each image token from different image areas receives heterogeneous computation allocation. Additionally, the method inherently achieves perfect load balance, as each expert processes an equal number of image tokens, which optimizes both efficiency and performance for the diffusion process.
@@
Recent advancements have leveraged MoEs to scale up dense diffusion models for image generation tasks. Several studies have focused on class-conditioned image generation. DTR (Park et al., 2023) and Switch-DiT (Park et al., 2024) approach the routing problem as multitask learning and utilize token-choice MoE for different denoising stages. Similarly, MEME (Lee et al., 2023) employs an ensemble of denoising experts specialized in distinct timestep ranges. Furthermore, some models have extended this approach for text-to-image generation. ERNIE-ViLG 2.0 (Feng et al., 2022) and eDiff-I (Balaji et al., 2022) employ ensembles of specialized experts for different denoising stages. Using U-Net as their backbone, these methods scale up to 24 billion parameters. However, these models activate only one expert at each timestep, which reduces efficiency at large scales. Recently, some works have shifted towards sparse MoEs. SegMoE (Puigcerver et al., 2023) employs top-k routing to scale up SDXL (Podell et al., 2023) with up to four experts. RAPHAEL (Xue et al., 2023) combines the concepts of distinct timestep experts and token-choice MoEs. DiT-MoE (Fei et al., 2024) applies token-choice routing to produce a sparse DiT with up to 16 billion parameters. Nevertheless, these methods often suffer from load imbalance issues inherent to the token-choice scheme, requiring an additional auxiliary balancing loss to mitigate this problem. The most similar work to ours is MoMa (Lin et al., 2024), which also utilizes expert-choice routing. However, MoMa explores model scaling only on a relatively small scale, with its largest model reaching 7 billion parameters. In contrast, we leveraged the proposed adaptive routing to scale EC-DiT up to 97 billion parameters and demonstrated the DiT scaling potential on a much larger scope.
5 Conclusion and Future Work
@@ Acknowledgments
We would like to thank John Peebles, Xianzhi Du, Chen Chen, Wenze Hu, Riley Roberts, Wei Liu, Zhengfeng Lai, Vasileios Saveris, Peter Grasch, and Yinfei Yang for constructive feedbacks, data curations, and discussions.
Appendix A Detailed DSG Scores
@@
Table 3 presents the detailed DSG scores for Dense and EC-DiT across various model configurations. Scaling with EC-DiT consistently improves performance in several categories, such as Counting and Text, and leads to better overall performance.
Appendix B GenEval Comparison with Token-choice Baseline
@@
EC-XXL-8E 69.11 99.84 85.43 68.59 81.85 20.31 58.61
EC-XXL-16E 69.43 99.34 83.46 71.62 83.83 20.94 57.42
EC-XXL-32E 70.86 99.53 86.47 72.34 84.39 21.00 61.39
Table 4 further compares the GenEval performance of EC-DiT with the token-choice baselines (Lepikhin et al., 2020a). By leveraging global image information and adaptive computation, EC-DiT demonstrates superior performance over the token-choice baseline across most evaluation categories.
Appendix C More Generated Samples
Figure 9: Generated samples from the Partiprompts subset (Yu et al., 2022). For each prompt, the model generates four candidates, and the one with the best visual satisfaction is selected.

### v5

@@ 3.5 Comparison with Token-choice Routing
Figure 5 compares EC-DiT with models scaled using token-choice MoE (Yatharth Gupta, 2024; Xue et al., 2023; Fei et al., 2024). For a fair comparison, we maintained identical experimental settings, except for replacing the routing strategy with token-choice and incorporating an auxiliary load-balancing loss in the training objective. The token-choice baseline uses top-2 routing, which matches the activation size of EC-DiT with a capacity factor of $f_{c}=2.0$. Our method consistently demonstrates superior training convergence and performance throughout the entire training period. Notably, EC-DiT with 8 experts rivals the token-choice baseline with 16 experts in both generation quality and text-image alignment, while EC-DiT with more experts significantly outperforms this token-choice baseline. One reason for the slower training convergence in token-choice routing is the equal compute allocation to different image areas, which makes the compute less adaptive. Detailed image areas may not receive sufficient computation to render complex patterns, while simpler areas waste the computation they receive. Moreover, token-choice routing processes each token independently without considering sequence information, leading to a lack of perception of the overall relationships between local tokens. Additionally, the inherent load imbalance in token-choice routing results in some experts being overloaded with an excessive number of tokens and creating a training bottleneck (Zhou et al., 2022). Appendix C provides a comprehensive comparison of EC-DiT and the token-choice baseline regarding GenEval performance. As shown in Table 5, EC-DiT outperforms the token-choice model across all tested configurations, which further validates the advantages of EC-DiT in text-image alignment.
@@ 4 Related Work
Many MoE models utilize a token-choice routing strategy, where each token independently selects the top-$k$ experts for processing. Sparsely Gated MoE (Shazeer et al., 2017) first introduced top-$k$ gating in LSTM models. Such routing strategy was later extended to transformers for language modeling in GShard (Lepikhin et al., 2020), Switch Transformer (Fedus et al., 2021), GLaM (Du et al., 2022), DeepSpeed-MoE (deepspeed), and ST-MoE (Zoph et al., 2022). To address potential load imbalances among experts, these token-choice approaches typically incorporate an auxiliary load-balancing loss (Shazeer et al., 2017). Alternatively, Zhou et al. (2022) proposed an expert-choice routing in text domain that selects the most suitable text tokens for each expert. Inspired by this, we adapt and scale up the DiT model using expert-choice routing tailored specifically for diffusion-based generative tasks. With this approach, each image token from different image areas receives heterogeneous computation allocation. Additionally, the method inherently achieves perfect load balance, as each expert processes an equal number of image tokens, which optimizes both efficiency and performance for the diffusion process.
@@
Recent advancements have leveraged MoEs to scale up dense diffusion models for image generation tasks. Several studie have focused on class-conditioned image generation. DTR (Park et al., 2023) and Switch-DiT (Park et al., 2024) approach the routing problem as multitask learning and utilize token-choice MoE for different denoising stages. Similarly, MEME (Lee et al., 2023) employs an ensemble of denoising experts specialized in distinct timestep ranges. Furthermore, some models have extended this approach for text-to-image generation. ERNIE-ViLG 2.0 (Feng et al., 2022) and eDiff-I (Balaji et al., 2022) employ ensembles of specialized experts for different denoising stages. Using U-Net as their backbone, these methods scale up to 24 billion parameters. However, these models activate only one expert at each timestep, which reduces efficiency at large scales. Recently, some works have shifted towards sparse MoEs. SegMoE (Puigcerver et al., 2023) employs top-k routing to scale up SDXL (Podell et al., 2023) with up to four experts. RAPHAEL (Xue et al., 2023) combines the concepts of distinct timestep experts and token-choice MoEs. DiT-MoE (Fei et al., 2024) applies token-choice routing to produce a sparse DiT with up to 16 billion parameters. Nevertheless, these methods often suffer from load imbalance issues inherent to the token-choice scheme, requiring an additional auxiliary balancing loss to mitigate this problem. The most similar work to ours is MoMa (Lin et al., 2024), which also utilizes expert-choice routing. However, MoMa explores model scaling only on a relatively small scale, with its largest model reaching 7 billion parameters. In contrast, we leveraged the proposed adaptive routing to scale EC-DiT up to 97 billion parameters and demonstrated the DiT scaling potential on a much larger scope.
5 Conclusion
@@ 6 Ethics Statement
7 Reproducibility Statement
In Section 2, we present our method with detailed formulations, examples, and visual demonstrations to clarify the model structure and mechanism (such as Figure 2). In Section 3, we introduce the training dataset and describe the most effective model parameters and components. We also provide specific details regarding the hardware used for training and inference like Figure 5. Additionally, we report the training dynamics of the model, depicted in Figure 8. Through various tables and figures, we present the performance of the proposed method from multiple perspectives, such as Table 2, Figure 3, and Figure 7. We compared our approach with existing methods targeting similar domains or overlapping methodologies, including both theoretical analysis and experimental evaluations (e.g., Figure 5). Notably, all evaluation experiments were conducted on public datasets or benchmarks, such as MSCOCO, GenEval, FID, and CLIP Score. To further ensure reproducibility, we plan to release the model weights contingent on the acceptance of this work.
@@ Acknowledgments
We would like to thank John Peebles, Xianzhi Du, Chen Chen, Wenze Hu, Riley Roberts, Wei Liu, Zhengfeng Lai, Vasileios Saveris, Peter Grasch, and Yinfei Yang for constructive feedback and discussions.
Appendix A Future Work
Additionally, the adaptive routing mechanism of EC-DiT can seamlessly integrate into any sequence- or chunk-wise encoding and generation process. For instance, adaptive routing can be directly applied to the tokenization (Gafni et al., 2022) commonly used in mixed-modal, early-fusion language models (Team, 2024). Similarly, in multi-modal models unifying autoregressive and diffusion paradigms (Zhou et al., 2024), expert-choice routing can be naturally integrated into the diffusion-based image generation component. Furthermore, EC-DiT could potentially enhance long-video generation methods (Chen et al., 2023b; Guo et al., 2023), where the global context within each video chunk could be effectively utilized. We will also explore these directions in our future work.
Appendix B Detailed DSG Scores
@@
Table 4: DSG comparison with token-choice baselines. GShard (GS) are implemented with two base model configurations (XL and XXL). GShard uses top-2 token-choice routing.
GS-XL-8E 73.47 73.27 51.43 63.73 86.00 73.50 86.70 78.67 76.37
GS-XL-16E 73.90 74.30 52.80 62.80 86.90 74.00 85.90 81.00 74.30
GS-XL-32E 74.07 74.13 52.17 64.53 86.50 75.23 86.60 78.47 77.23
GS-XXL-8E 75.10 76.10 53.80 68.60 87.40 72.70 87.70 80.50 76.00
GS-XXL-16E 75.30 74.50 52.83 66.57 87.27 77.67 87.73 80.63 78.33
GS-XXL-32E 75.40 76.40 52.70 67.20 87.10 78.90 88.40 79.80 76.30
Table 3 presents the detailed DSG scores for Dense and EC-DiT across various model configurations. Scaling with EC-DiT consistently improves performance in several categories, such as Counting and Text, and leads to better overall performance. Table 4 shows that EC-DiT demonstrates superior performance over the token-choice baselines across most evaluation DSG categories.
Appendix C GenEval Comparison with Token-choice Baseline
@@
EC-DiT–XXL-8E 69.11 99.84 85.43 68.59 81.85 20.31 58.61
EC-DiT–XXL-16E 69.43 99.34 83.46 71.62 83.83 20.94 57.42
EC-DiT–XXL-32E 70.86 99.53 86.47 72.34 84.39 21.00 61.39
Table 5 further compares the GenEval performance of EC-DiT with the token-choice baselines (Lepikhin et al., 2020). By leveraging global image information and adaptive computation, EC-DiT demonstrates superior performance over the token-choice baseline across most evaluation categories.
Appendix D Calculating Activated Parameters Increment
We present the pseudocode of calculating the activated parameters increment in Algorithm 2. Here, the parameters hidden_dim, num_sparse_layers, num_heads, and num_kv_heads for each model configuration are specified in Table 1. We set ffn_factor=4.0 and capacity_factor=2.0 globally.
Algorithm 2 Pseudocode for Activated Parameters Increment Calculation
# Dense FFN Size
dense_ffn_size = 2 * hidden_dim * hidden_dim * ffn_factor
# Router Size
router_size = num_experts * hidden_dim
# Attention Size
attn_size = hidden_dim * (num_heads + num_kv_heads * 2 + num_heads) * attn_key_dim
# Activated Dense Total
activated_dense_total = (attn_size + dense_ffn_size) * num_sparse_layers
# Activated Sparse Total
activated_sparse_total = (
attn_size
+ 2 * hidden_dim * hidden_dim * ffn_factor * capacity_factor
+ router_size
) * num_sparse_layers
# Activated Total Increment
activated_increment = activated_sparse_total - activated_dense_total
Appendix E More Generated Samples and Visual Comparisons
Figure 9: Visual comparison of three model configurations on the Partiprompts subset. For each prompt, the model generates four candidates, and the one with the best visual satisfaction is selected.
We provide an additional comparison on the PartiPrompt subset (Yu et al., 2022) across different model configurations and also benchmark against two external text-to-image generative models: SD3-Large (Esser et al., 2024) and FLUX.1[dev] (Black Forest Labs, 2024).
Both SD3-Large and FLUX.1[dev] have comparable or higher activation sizes than EC-DiT-M-64E. Despite this, EC-DiT demonstrates superior text-to-image alignment compared to these models. For example, in Prompt #3, FLUX.1[dev] fails to align with the textual elements holding a cane and holding a garbage bag simultaneously, nor adhering to the requested style of abstract cubism. In contrast, EC-DiT effectively captures both the described actions and the specified artistic style. Similarly, in Prompt #5, SD3-Large and FLUX.1[dev] fail to include the background detail of Dois Irmãos in Rio de Janeiro, while EC-DiT correctly incorporates the described view.
Furthermore, EC-DiT generates images with lower hallucination. For instance, in Prompt #2, SD3-Large hallucinates a third claw on the hamster, and FLUX.1[dev] adds an unnecessary extra "!" mark. In contrast, EC-DiT ’s generation appears visually accurate and reasonable.
Figure 10: Visual comparison of EC-DiT-M-64E and two external text-to-image models on the Partiprompts subset. Numbers in parentheses indicate the activated parameter size for each model. All models leverage rectified flow for generation. SD3-Large inherently generates at $1024\times 1024$, while FLUX.1[dev] generates at $512\times 512$.
Figure 11: More generated samples with EC-DiT-M-64E on the Partiprompts subset. For each prompt, the model generates four candidates, and the one with the best visual satisfaction is selected.